OpenStudy (anonymous):

If the car can brake with an acceleration of -14.5m/s^2 howling will it take to come to a complete stop from a speed of 75m/s

OpenStudy (anonymous):

use v=u+at

OpenStudy (anonymous):

What does the u stand for? @surjithaye

OpenStudy (anonymous):

you followed or any problen m. u=75 m/s

OpenStudy (anonymous):

u=initial velocity.

OpenStudy (anonymous):

u=75m/s final speed v=0m/s a=-15m/s^2 plug the values and find the value of t.

OpenStudy (anonymous):

correction a=-14.5m/s^2